OpenStudy (anonymous):

a tennis player serves a ball horizontally giving it a speed of 24m/s from a height of 2.5m. the player is 12m from the net. the top of the net is 0.90m above the court surface. the ball clears the net and lands on the other side. air resistance is negligible, by what distance does the ball clear the net?

OpenStudy (callisto):



OpenStudy (anonymous):

Thanks for the diagram, lol I drew mine like that too

OpenStudy (callisto):

Consider the horizontal motion: \[s_x = u_xt\]\[12 = 24t\]\[t=0.5\] Consider the vertical motion: \[s_y = ut - \frac{1}{2}gt^2\]\[s_y = -\frac{1}{2}g(0.5)^2=...\]

OpenStudy (callisto):

Distance required = (2.5-\(s_y\)) - 0.9 = ...